Zhongmin is 50 years old. His son, Mingyong is 24 years old. How many years ago was Zhongmin three times as old as Mingyong?
Answer by ewatrrr(24785)

 
Hi,
Let x represent # of years ago that Zhongmin was three times as old as Mingyong
Question states***Zhongmin is 50 years old. His son, Mingyong is 24 years old
50-x = 3(24-x)
2x = 22
x = 11yr, # of years ago that Zhongmin was three times as old as Mingyong
CHECKING our Answer***
(50-11) = 39 = 3*(24-11) = 39